OpenStudy (anonymous):

What is the value of f[g(6)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

g(6)=2(6)-5=7 f[g(6)]=2(7)+1=15 i think it is like that